Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 543
Entry time: 09 Jul 2018, 14:32
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 13:00 |
|
|
SPB
- Mirrors installatiostarting now. Preliminary motion tests conducted last friday were successful. Still closed loop for encoders signal has to be implemented. AE will work on that.
- AGIPD firmware updated. Remaining tests ongoing. New Karabo release will be deployed tomorrow.
- Peltiers being tested.
SQS
- End to end tests procceding. Laser incoupling and NQS chambers being set up.
- Loop 5 needed for next week.
- M pot firmware issue is being adreesed by CAS
- It is not possible to switch the ion pumps on and off remotely using Karabo. AE and CAS to look into that.
AE
- SASE3 exp. Loops preparation ongoing
- SPB Loop 4 update being prepared
- SASE3 tunnels interlock definitions are being updated for some diagnostic component
- SASE2 MID CRLs being tested.
CAS
- SPB 2.2.4.1 Karabo update scheduled for tomorrow morning between 7 and 9.
|